OpenStudy (anonymous):

Find the Maclaurin series for f(x) using the definition of a Maclaurin series. [Assume that f has a power series expansion. Do not show that Rn(x) → 0.] f(x) = 4 sinh 4x f(x)= sum n=0 to infinity ________?

OpenStudy (anonymous):

i found (x^(2n+1)/((2n+1)factorial)) ????????

OpenStudy (perl):

can you define sinh x

OpenStudy (anonymous):

0,1,0,1,0 ???

OpenStudy (perl):

no i mean, in terms of e^x

OpenStudy (perl):

whats the derivative of sinh x , cosh x ?

OpenStudy (perl):

ok sinh and cosh have nice property

OpenStudy (perl):

sinh x ' = cosh x cosh x ' = sinh x

OpenStudy (anonymous):

so is my answer correct:S

OpenStudy (perl):

OS crashed :(

OpenStudy (perl):

f(x) = f(0) + f ' (0) x + f '' (0)/2! * x^2 + f ' ' ' (0) / 3! + ... f(x) = 4sinh(4x) f'(x) = 4 cosh (4x)*4 = 16 cosh (4x) f '' (x) = 16 sinh(4x) * 4 = 64 sinh(4x) f'''(x) = 256 cosh(x) f''''(x) = 1024 sinh (x) f(0) = 0 f'(0)= 16 f ''(0) = 0 f '''(0) = 256

OpenStudy (perl):

a much simpler solution is , we know the Maclaurin series for sinh x :